OpenStudy (cenaida):

If the length of a large banner is 4x + 5 and the width is half as long, what is the area? A. 6x² + 20x + 12.5 B. 8x² + 17x + 12.5 C. 8x² + 20x + 12.5 D. 8x² + 22x + 12.5

OpenStudy (anonymous):

2x+2.5 multiplied by 4x+5

OpenStudy (shamil98):

area = l x w length = 4x +5 width = (4x+5)/2 (4x + 5 )(2x + 2.5)

OpenStudy (anonymous):

The area of a rectangle is it's length time its width. \[A = ( length)(width)\] The width is the same as half the length. \[A = (length) \left( \frac{ length }{ 2 } \right)\] The length is equal to 4x+5. \[A = (4x+5)\left( \frac{ 4x+5 }{ 2 } \right)\] You can put the half out of the operation. \[A=\frac{ (4x+5)(4x+5) }{ 2 }=\frac{ (4x+5)^2 }{ 2 }\]

OpenStudy (anonymous):

\[\frac{ (4x+5)^{2} }{ 2 }= \frac{ 16x^{2} + 40x + 25 }{ 2 }=8x^{2} + 20 x + 12.5\]

OpenStudy (anonymous):

And. That's the area.

OpenStudy (cenaida):

thanks for your great explanation everyone @Lessis